Marvin wants to purchase a rectangular rug. it has an area of 80 square ft. he cannot remember the length and width, but he remembers that the length was 8 more than some number and the width was 8 more than the same number.
What is the quadratic eqation that can be made from the above information?

Answer by oberobic(2304)

Let 'some number' be 'x'.
L = length = x + 8
W = width = x + 8
.
The transitive property teaches that if a=b and b=c, then a=c.
Therefore, L = W.
That means the rug is square.
Of course, a square is an equilateral rectangle, so that is okay.
.
A = area = L*W = L^2 = W^2 = 80 sq. ft.
.
L = sqrt(80) = sqrt(16*5) = sqrt(16)*sqrt(5) = 4*sqrt(5) = 8.94427191
.
So L = W = 8.94427191
.
Returning to the setup, L = x+8 = 8 + .94427191.
So x = .94427191
No wonder he could not remember it.
.
Done.